# Broadening of the Divertor Heat Flux Profile in High Confinement Tokamak Fusion Plasmas with Edge Pedestals Limited by Turbulence in DIII-D

### v1

@@ Abstract
The XGC simulations of toroidal mode number N>5N\!>5\! modes show overall consistency with measurements of the changes in shorter wavelength fluctuation intensity. Fig. 8(a) shows $\delta n/n$ for the dominant toroidal mode number ($N$) fluctuations propagating in the electron direction, for WPQH-mode ($N=7$) and QH-Mode ($N=9$), indicating that XGC simulations find a relative increase for this toroidal mode similar to that measured in Figs. 4(a) in the steep gradient pedestal region. We have implemented synthetic BES in XGC to attempt validation of the XGC results against experimental observations at higher frequencies corresponding to $N>5$; preliminary results nearly match the total measured BES $\delta n/n$ profile for the QH case. The corresponding cross-phase between density and potential fluctuations $(\delta n,\delta\phi)$ from XGC is shown in Fig. 8(b). Noting that an “adiabatic” or Boltzmann electron density response would be in phase with the electrostatic potential, a strong non-adiabatic electron behavior (typical of the trapped electron response) is indicated by the non-zero cross-phase, peaking near $\psi_{N}=0.96$ for $N=9$ in the QH case and around $\psi_{N}=0.98$ for $N=7$ in the WPQH case, corresponding to the peaks in the electron directed turbulence $\delta n/n$ in Fig. 8(a). The non-zero cross-phase, typical of TEM turbulence, produces significant electron particle and thermal transport. From XGC, the lab frame frequency of the QH feature is 330 kHz, while the poloidal wavenumber $k_{\theta}=0.63$ cm-1, consistent with the measurements in Fig. 4(c, d). In the plasma frame moving with the $E\times B$} velocity, the WPQH-Mode case shows both ion and electron modes spanning a range of wavenumbers $k_{\theta}\sim 0.1-0.6$ cm-1 (Fig. 8(c)), while the QH-Mode case shows mainly electron modes with $k_{\theta}\sim 0.4-0.7$ cm-1 (Fig. 8(d)). Similar results are seen inward of this radius, while very close to the separatrix, the electron directed turbulence moves to longer wavelengths, $k_{\theta}\sim 0.1-0.4$ cm-1, but remains stronger in the QH case than in the WPQH case.
Finally, scaling arguments based on theory and simulation suggest that in future machines, pedestal turbulent transport may be sufficiently limiting to naturally maintain stability to edge localized modes (ELMs) [ernst:2018, ernst:2022aps]. Turbulent transport reduction by sheared flows is predicted to scale with the normalized ion gyroradius $\rho_{*}=\rho_{i}/a$ where $a$ is the plasma minor radius [kotschenreuther:1996]. The radial electric field $E_{r}$ in the edge pedestal is $Zen_{i}E_{r}\sim dp_{i}/dr\sim p_{i}/\Delta$, where $Zen_{i}$ is the ion charge density, $p_{i}=n_{i}T_{i}$ is the ion pressure with $T_{i}$ the ion temperature, $r$ is the minor radius, and $\Delta$ is the gradient scale length The $E\times B$} shear rate for turbulence is then $\gamma_{E}\sim B^{-1}dE_{r}/dr\sim E_{r}/(B\Delta)$ where $B$ is the magnetic field. Estimating the growth rate for ion scale drift-type instabilities as $\gamma_{\mathrm{lin}}\sim v_{\mathrm{thi}}/\Delta$, where $v_{\mathrm{thi}}=(2T_{i}/m_{i})^{1/2}$ is the ion thermal speed, the reduction of turbulent transport by $E\times B$} shear will depend on a parameter which scales as $\gamma_{E}/\gamma_{\mathrm{lin}}\sim\rho_{i}/\Delta\sim(a/\Delta)\rho_{*}$. The ratio of pedestal turbulent heat flux $Q$ to the gyroBohm flux $Q_{\mathrm{gB}}$ is predicted to increase asymptotically as a result of the decreased $E\times B$} shear at low $\rho_{*}$ in future machines according to [zhang:1992, kotschenreuther:2017, hatch:2018]
${Q_{\hskip 6.02777pt}\over Q_{\mathrm{gB}}}\sim\left({\gamma_{E}\over\gamma_{\mathrm{lin}}}\right)^{-2}\sim\left({\Delta\over a}\right)^{2}{1\over\rho_{*}^{2}}\sim{\beta_{p}^{2{\alpha_{1}}}\over\rho_{*}^{2}}$
where $\beta_{p}=8\pi p/B_{p}^{2}$ is the poloidal beta parameter, which we have included by invoking the EPED scaling for the pedestal width as limited by the onset of kinetic ballooning modes, $\Delta\sim\beta_{p}^{\alpha_{1}}$, where $\alpha_{1}\sim 0.5-0.75$ [snyder:2011, snyder:2012]. Assuming an H-Mode pedestal is formed, pedestal turbulent transport could be approximately an order of magnitude stronger in future high magnetic field machines as a result of the factor of three reduction in $\rho_{*}$, assuming the pedestal width follows EPED scaling (KBM onset) or a “softer” version of it with $\alpha_{1}<0.75$ due to other less virulent instabilities. Intrinsically Non-ELMing operating regimes such as Wide Pedestal Quiescent H-Mode and Quiescent H-Mode offer a high-performance solution without ELMs at relevant low pedestal top collisionalities, with the added benefit of broadened divertor heat flux profiles when the pedestal turbulent transport is sufficiently strong.
@@ Appendix A Appendix
The integral heat flux profile width $\lambda_{\mathrm{int}}$ can be related to the Eich fit parameters $(\lambda_{q},S)$ by $\lambda_{\mathrm{int}}=\int(q_{\|}(s)-q_{\mathrm{BG}})ds/(q_{\|0})\approx\lambda_{q}+1.64\,S$ [makowski:2012]. The Eich function is given by
@@
where $\bar{s}=s-s_{0}=(R_{\mathrm{mid}}-R_{\mathrm{sep}})f_{x}$ is the midplane radial coordinate and $f_{x}=R_{\mathrm{div}}B_{\mathrm{pol}}^{\mathrm{div}}/R_{\mathrm{mp}}B_{\mathrm{pol}}^{\mathrm{mp}}\approx 5.3$ (for our cases) is the flux expansion effective area factor from midplane to divertor. The two parameters $(\lambda_{q},S)$ characterize the midplane exponential decay length and the Gaussian width, where $S$ represents competition between parallel and perpendicular heat transport between the X-point and strike point [eich:2013]. The Eich function is the convolution of an exponential decay in the scrape-off layer and a Gaussian diffusive width characterized by S. Note the IR data for these experiments has insufficient coverage to constrain S, which is obtained from Langmuir probe measurements.
Profile fits, shown in Fig. A.1, are constrained by Langmuir probe measurements of separatrix electron temperature $T_{e}^{\mathrm{s}ep}$ and pedestal carbon and deuterium charge exchange measurements, which show constant carbon concentration across the pedestal. Equilibrium reconstructions of discharge #184833 in WPQH and QH phases, including the bootstrap current and realistic heating profiles from interpretative transport simulations (TRANSP) are used as input to the XGC code.
@@
Using the standard sheath heat transmission coefficient $\gamma_{\mathrm{SH}}=7$ [Stangeby:2000] to relate total parallel heat flux from electrons and ions to Langmuir probe measured saturation current ($J_{\mathrm{sat}}$) and electron temperature ($T_{e}$) via $q_{\|}^{\mathrm{tot}}=\gamma_{\mathrm{SH}}J_{\mathrm{sat}}T_{e}/e$, close agreement is obtained between Langmuir probe measured divertor heat flux profiles and IR thermography measured heat flux profiles, as shown in Fig. A.3. The IR measurements are shown for corresponding cases where available. This agreement translates to agreement within statistical measurement uncertainty in the $\lambda_{q}$ values inferred from Langmuir probes and IR, as shown in Fig. A.2. Importantly, the $\lambda_{q}$ inferred from Langmuir probes is independent of the assumed $\gamma_{\mathrm{SH}}$. Because Langmuir probes measure mainly electrons, this result can be viewed as evidence that the electron contribution to the divertor heat flux underlies the broadening of $\lambda_{q}$ measured by IR.

### v2

@@ Abstract
The XGC simulations of toroidal mode number N>5N\!>5\! modes show overall consistency with the fluctuation measurements described above. Fig. 8(a) shows $\delta n/n$ for the dominant toroidal mode number ($N$) fluctuations propagating in the electron direction, for WPQH-mode ($N=7$) and QH-Mode ($N=9$), indicating that XGC simulations find a relative increase for this toroidal mode similar to that measured in Figs. 4(a) in the steep gradient pedestal region. We have implemented synthetic BES in XGC to attempt validation of the XGC results against experimental observations at higher frequencies corresponding to $N>5$; preliminary results nearly match the total measured BES $\delta n/n$ profile for the QH case. The corresponding cross-phase between density and potential fluctuations $(\delta n,\delta\phi)$ from XGC is shown in Fig. 8(b). Noting that an “adiabatic” or Boltzmann electron density response would be in phase with the electrostatic potential, a strong non-adiabatic electron behavior (typical of the trapped electron response) is indicated by the non-zero cross-phase. This produces significant electron particle and thermal transport, typical of TEM turbulence. From XGC, the lab frame frequency of the QH feature is 330 kHz, while the poloidal wavenumber $k_{\theta}=0.63$ cm-1, consistent with the measurements in Fig. 4(c, d). In the plasma frame moving with the $E\times B$} velocity, the WPQH-Mode case shows both ion and electron modes spanning a range of wavenumbers, $k_{\theta}\sim 0.1-0.6$ cm-1 (Fig. 8(c)), while in contrast, the QH-Mode case shows mainly electron modes with $k_{\theta}\sim 0.4-0.7$ cm-1 (Fig. 8(d)).
Finally, scaling arguments based on theory and simulation suggest that in future machines, pedestal turbulent transport may be sufficiently limiting to naturally maintain stability to edge localized modes (ELMs) [ernst:2018, ernst:2022aps] so that Non-ELMing regimes naturally arise (see Appendix for details). Intrinsically Non-ELMing operating regimes such as Wide Pedestal Quiescent H-Mode and Quiescent H-Mode offer a high-performance solution without ELMs at relevant low pedestal top collisionalities, with the added benefit of broadened divertor heat flux profiles when the pedestal turbulent transport is sufficiently strong.
@@ Appendix A Appendix
Eich Function. The integral heat flux profile width $\lambda_{\mathrm{int}}$ can be related to the Eich fit parameters $(\lambda_{q},S)$ by $\lambda_{\mathrm{int}}=\int(q_{\|}(s)-q_{\mathrm{BG}})ds/(q_{\|0})\approx\lambda_{q}+1.64\,S$ [makowski:2012]. The Eich function is given by
@@
where $\bar{s}=s-s_{0}=(R_{\mathrm{mp}}-R_{\mathrm{LCFS}})f_{x}$ is the radial coordinate at the divertor, referring to the departure in midplane major radius from the LCFS, and $f_{x}=R_{\mathrm{div}}B_{\mathrm{pol}}^{\mathrm{div}}/R_{\mathrm{mp}}B_{\mathrm{pol}}^{\mathrm{mp}}\approx 5.3$ (for our cases) is the flux expansion effective area factor from midplane to divertor. The two parameters $(\lambda_{q},S)$ characterize the midplane exponential decay length and the Gaussian width, referred to the midplane, where $S$ represents competition between parallel and perpendicular heat transport between the X-point and strike point [eich:2013]. The Eich function is the convolution of an exponential decay in the scrape-off layer and a Gaussian diffusive width characterized by S. Note the IR data for these experiments has insufficient coverage to constrain S, which is obtained from Langmuir probe measurements.
@@
Profile Fits. Pedestal profiles, shown in Fig. A.1, are constrained by Langmuir probe measurements of separatrix electron temperature $T_{e}^{\mathrm{s}ep}$ and pedestal carbon and deuterium charge exchange measurements, which show constant carbon concentration across the pedestal. Equilibrium reconstructions of discharge #184833 in WPQH and QH phases, including the bootstrap current and realistic heating profiles from interpretative transport simulations (TRANSP) are used as input to the XGC code.
Divertor Heat Flux Measurements. Using the standard sheath heat transmission coefficient $\gamma_{\mathrm{SH}}=7$ [Stangeby:2000] to relate total parallel heat flux from electrons and ions to Langmuir probe measured saturation current ($J_{\mathrm{sat}}$) and electron temperature ($T_{e}$) via $q_{\|}^{\mathrm{tot}}=\gamma_{\mathrm{SH}}J_{\mathrm{sat}}T_{e}/e$, close agreement is obtained between Langmuir probe measured divertor heat flux profiles and IR thermography measured heat flux profiles, as shown in Fig. A.3. The IR measurements are shown for corresponding cases where available. This agreement translates to agreement within statistical measurement uncertainty in the $\lambda_{q}$ values inferred from Langmuir probes and IR, as shown in Fig. A.2. Importantly, the $\lambda_{q}$ inferred from Langmuir probes is independent of the assumed $\gamma_{\mathrm{SH}}$. Because Langmuir probes measure mainly electrons, this result can be viewed as evidence that the electron contribution to the divertor heat flux underlies the broadening of $\lambda_{q}$ measured by IR.
@@
Scaling Pedestal Turbulent Transport to Future Machines. Here we provide details on the scaling of pedestal turbulent transport with ion gyroradius, which varies inversely with magnetic field. Turbulent transport reduction by sheared flows is predicted to scale with the normalized ion gyroradius, $\rho_{*}=\rho_{i}/a$, where $a$ is the plasma minor radius [kotschenreuther:1996]. In future machines, $\rho_{*}$ is expected to be approximately three times smaller than in present tokamaks. The radial electric field $E_{r}$ in the edge pedestal is $Zen_{i}E_{r}\sim dp_{i}/dr\sim p_{i}/\Delta$, where $Zen_{i}$ is the ion charge density, $p_{i}=n_{i}T_{i}$ is the ion pressure with $T_{i}$ the ion temperature, $r$ is the minor radius, and $\Delta$ is the gradient scale length The $E\times B$} shear rate for turbulence is then $\gamma_{E}\sim B^{-1}dE_{r}/dr\sim E_{r}/(B\Delta)$ where $B$ is the magnetic field. Estimating the growth rate for ion scale drift-type instabilities as $\gamma_{\mathrm{lin}}\sim v_{\mathrm{thi}}/\Delta$, where $v_{\mathrm{thi}}=(2T_{i}/m_{i})^{1/2}$ is the ion thermal speed, the reduction of turbulent transport by $E\times B$} shear will depend on a parameter which scales as $\gamma_{E}/\gamma_{\mathrm{lin}}\sim\rho_{i}/\Delta\sim(a/\Delta)\rho_{*}$. The ratio of pedestal turbulent heat flux $Q$ to the gyroBohm flux $Q_{\mathrm{gB}}$ is predicted to increase asymptotically as a result of the decreased $E\times B$} shear at low $\rho_{*}$ according to [zhang:1992, kotschenreuther:2017, hatch:2018]
${Q_{\hskip 6.02777pt}\over Q_{\mathrm{gB}}}\sim\left({\gamma_{E}\over\gamma_{\mathrm{lin}}}\right)^{-2}\sim\left({\Delta\over a}\right)^{2}{1\over\rho_{*}^{2}}\sim{\beta_{p}^{2{\alpha_{1}}}\over\rho_{*}^{2}}.$
Here $\beta_{p}=8\pi p/B_{p}^{2}$ is the poloidal beta parameter, which we have included by invoking the EPED scaling for the pedestal width as limited by the onset of kinetic ballooning modes, $\Delta\sim\beta_{p}^{\alpha_{1}}$, where $\alpha_{1}\sim 0.5-0.75$ [snyder:2011, snyder:2012]. Assuming an H-Mode pedestal is formed, pedestal turbulent transport could be approximately an order of magnitude stronger in future high magnetic field machines as a result of the factor of three reduction in $\rho_{*}$, assuming the pedestal width follows EPED scaling (KBM onset) or a “softer” version of it with $\alpha_{1}<0.75$ due to other less virulent instabilities.